OpenStudy (anonymous):

Which equation contains the points (5, 3) and (2, 6)? A. y = x + 2 B. y = x - 2 C. y = - x + 8 D. y = - x - 8

sam (.sam.):

Well you can find the slope first, do you know how to find it?

sam (.sam.):

\[\Large (x_1,y_1),(x_2,y_2)\] \[\Large Slope=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Yes.

sam (.sam.):

So what do you get?

sam (.sam.):

Hmm maybe try substituting the values \[\Large Slope=\frac{6-3}{2-5}\]

sam (.sam.):

We get \[Slope=-1\]

sam (.sam.):

Then use the point slope formula to create a new equation,

sam (.sam.):

\[y-y_3=Slope(x-x_4) \\ \\ y-3=-(x-5)\] Then simplify it

OpenStudy (anonymous):

How'd you get x4?

sam (.sam.):

The x4 is just another variable

sam (.sam.):

We used y3 and x4 from any one of the coordinates, (5, 3) and (2, 6)

sam (.sam.):

I used (5,3)

OpenStudy (anonymous):

I got y = x - 2

sam (.sam.):

How do you get that? We got \[ y-3=-(x-5) \\ \\ y-3=-x+5 \\ \\ y=-x+5+3 \\ \\ y=-x+8\]

sam (.sam.):

@strugglingmathkid